OpenStudy (anonymous):

What is $\left(4\dfrac{5}{8}\right)^{55} \cdot \left(\dfrac{8}{37}\right)^{55}$?

OpenStudy (anonymous):

\[ \left(4\dfrac{5}{8}\right)^{55} \cdot \left(\dfrac{8}{37}\right)^{55}\]

OpenStudy (anonymous):

that's the question

OpenStudy (anonymous):

yeah i was just translating it

OpenStudy (anonymous):

the idea is that this is the same as \[\left(4\tfrac{5}{8}\times \frac{8}{37}\right)^{55}\]

OpenStudy (anonymous):

compute the number inside first maybe it will be easy

OpenStudy (anonymous):

lol let me know when you get \(1\)

OpenStudy (anonymous):

oh I got it